# CGCE [Uncaptioned image]: Classifier-Guided Concept Erasure in Generative Models

### v1

@@ 1 Introduction
Several strategies have been proposed to mitigate the generation of such undesirable content. A straightforward approach is dataset filtering [1], which involves removing unsafe images before training. However, this is often computationally expensive and impractical for every new concept because aggressive filtering can inadvertently degrade overall generative quality. Post-hoc solutions have emerged as more practical alternatives. Model fine-tuning approaches [8, 12, 40, 16, 30] permanently alter the model’s weights to erase or remap harmful concepts, but this process can be resource-intensive and often leads to a trade-off between erasure effectiveness and utility preservation of safe content. A distinct category is model editing [9, 10] which uses closed-form solutions to directly alter model weights without iterative training. While highly efficient, these edits can also be incomplete and may remain vulnerable to adversarial attacks that regenerate the erased concepts. Finally, training-free, filtering-based methods [28, 38] modify the generation process at inference time by steering the output away from undesired concepts. While flexible, these methods can be less robust and are often limited to specific model architectures or susceptible to being bypassed, especially when prompts express harmful concepts implicitly.
To address these limitations, we introduce Classifier-Guided Concept Erasure (CGCE), an efficient framework that offers a robust and versatile solution for safe visual generation, as shown in Fig. 1. A key strength of CGCE is its plug-and-play nature, allowing it to be applied to diverse generative models, such as modern T2I and T2V architectures, without altering the original model weights. This preserves the high-quality generative capabilities of the base model for safe content, a critical advantage over many fine-tuning methods that suffer from utility degradation. Our approach leverages a lightweight, specially trained classifier that operates in the text embedding space. This classifier is first trained on a synthetic dataset of paired safe and unsafe prompts, enabling it to accurately detect the presence of an undesired concept within a given text embedding. At inference time, our classifier serves a dual role. It first acts as a safeguard, inspecting the text embedding to detect the presence of an undesired concept. Upon detection, it then functions as a refiner, altering the embedding to neutralize the harmful semantics and guide it toward a safe semantic region. This process ensures a safe visual output from the T2I or T2V model while leaving safe prompts untouched, thereby preserving the model’s original generation quality. Furthermore, our framework is highly scalable and can simultaneously erase multiple distinct concepts by aggregating the guidance from several classifiers.
@@ 3 Proposed Method
We propose CGCE, a lightweight, plug-and-play framework for safeguarding T2I and T2V models without altering their original weights. First, we use a Large Language Model (LLM) to generate prompt pairs with and without a target concept to train a lightweight classifier. During inference, this classifier acts as a safeguard by detecting the undesired concept in the text embedding. Upon detection, it refines the embedding by steering it away from the concept’s semantic representation using gradient descent before it is passed to the generative model. This process is illustrated in Fig. 2.
@@ 3.1 Concept Classifier
Synthetic Data Generation. Our goal is to build a classifier capable of detecting target concepts within the text embedding space used by T2I and T2V models. These models typically process an input prompt $p$, through a text encoder like CLIP [24] or T5 [25] to generate a text embedding. To train our classifier, we first construct a dataset of prompt pairs using a LLM (e.g., Gemini [3], Qwen [31]). Each pair consists of a prompt containing the target concept (unsafe, denoted $p_{u}$) and a corresponding prompt that preserves similar content but excludes the concept (safe, denoted $p_{s}$). This paired data structure is crucial for isolating the specific semantic features of the target concept, which helps the classifier learn a more precise decision boundary and resolves ambiguity about which features to identify. To enhance the classifier’s robustness, we ensure the prompt dataset is diverse across various scenarios. For instance, when targeting the “nudity” concept, a pair would include an unsafe prompt, $p_{u}$: a photo of a nude girl” and its corresponding safe prompt, $p_{s}$: a photo of a girl”. The templates used for prompting the LLM are detailed in the Appendix for reproducibility.
@@
To create a single representation for the entire prompt, we first derive an importance score $\boldsymbol{s}\in\mathbb{R}^{n}$, by taking the maximum attention weight that each prompt token $i$ assigned to any of the concept tokens $j$ ($\boldsymbol{s}_{i}=\max_{j}\mathbf{A}_{ij}$). These scores are normalized via a softmax function to produce a final importance distribution $\boldsymbol{\alpha}=\text{softmax}(\boldsymbol{s})$. This distribution is used to compute an aggregated sequence representation $\boldsymbol{\varepsilon}_{seq}$, by taking a weighted sum of the attended features: $\boldsymbol{\varepsilon}_{seq}=\sum_{i=1}^{n}\boldsymbol{\alpha}_{i}\cdot\boldsymbol{\varepsilon}_{att,i}$. This final vector $\boldsymbol{\varepsilon}_{seq}$, which emphasizes the prompt’s concept-related features, is then passed through an MLP and a sigmoid function to yield the final probability $\boldsymbol{f}_{\theta}(\boldsymbol{\varepsilon}_{p},\boldsymbol{\varepsilon}_{c})$, indicating the likelihood that the prompt contains the target concept.
@@ 3.2 Concept Erasure with Classifier
Classifier as a Refiner. If the target concept is detected in an unsafe prompt $p_{u}$, the classifier then acts as a refiner to steer the embedding $\boldsymbol{\varepsilon}_{p_{u}}$ towards a region of the latent space that does not contain the concept. Intuitively, the objective is to find a refined embedding $\boldsymbol{\varepsilon}^{\prime}_{p_{u}}$ that minimizes the classifier’s prediction, effectively driving its output towards zero. We achieve this by iteratively updating the embedding using weighted gradient descent of the classifier’s output with respect to the embedding itself. The update rule at each refinement step $k$ is given by:
@@ 4 Experiments
ESD [8] $\times$ 8.59 16.70 30.15 56.84 9.20 73.94 15.98 30.42
MACE [20] $\times$ 6.02 8.17 8.09 5.61 4.10 66.20 15.14 29.33
UCE [9] $\times$ 6.23 10.27 18.01 14.39 7.90 73.24 17.74 30.30
RECE [10] $\times$ 6.77 20.66 28.68 12.98 36.10 72.53 16.24 30.95
DUO [22] $\times$ 12.03 24.24 23.90 31.23 28.60 82.39 15.83 31.28
STEREO [30] $\times$ 0.75 4.74 5.15 4.21 7.90 28.17 17.98 30.16
SLD-Medium [28] 26.64 73.16 70.96 97.54 52.30 52.11 20.38 30.71
SLD-Strong [28] 20.84 59.39 58.09 91.58 39.60 34.75 23.19 29.97
SLD-Max [28] 15.68 46.46 40.81 64.21 35.60 21.83 26.63 29.20
SAFREE [38] 12.35 39.18 48.90 74.74 44.70 80.99 19.55 30.69
@@ 4.1 Experimental Setup
Baselines. Following recent works [8, 9, 10, 20, 38, 30], we use SD-v1.4 [27] as our primary T2I model for comparison. We evaluate CGCE against eight baselines, including training-based methods (e.g., ESD [8], DUO [22], STEREO [30]), closed-form editing methods (e.g., UCE [9], MACE [20], RECE [10]), and training-free inference methods (e.g., SLD [28], SAFREE [38]). The evaluation spans three main concept removal tasks:
@@ 4.2 Main Results
Nudity Removal. As shown in Tab. 2 and Fig. 3, CGCE achieves state-of-the-art or highly competitive performance across all six red-teaming benchmarks for nudity removal, without requiring any modification to the model weights. Our approach ranks first on four of the six benchmarks (e.g., SixCD, P4D, RAB, and MMAD). On the I2P and UDA benchmarks, CGCE achieves the second-best ASR. Notably, CGCE offers a significantly better balance between robustness and utility preservation than the top-performing methods on these benchmarks. For example, while STEREO achieves a lower ASR on I2P, it shows a noticeable drop in utility on the COCO dataset compared to CGCE. Similarly, SLD-Max has a lower ASR on UDA but suffers from severe utility degradation. In contrast, CGCE maintains FID and CLIP scores that are almost identical to those of the original SD-v1.4 model, indicating that we achieve strong robustness with almost no loss in the model’s general generative ability. This is because our classifier can accurately identify safe and unsafe content, making no changes to safe samples. Furthermore, CGCE consistently and significantly outperforms SAFREE, another training-free approach that operates on text embeddings, across all benchmarks. This is because SAFREE’s reliance on token-level proximity to a predefined ”toxic subspace” is less effective against prompts where harmful concepts are expressed implicitly or contextually.
@@ 4.3 Generalization and Extensibility of CGCE
Generalization Across Diverse T2I Architectures. To demonstrate the versatility of CGCE, we extend our evaluation to a range of modern T2I backbones, including DiT-based models (e.g., SD-v3, FLUX.1-dev) and VAR models (e.g., Switti-AR, Infinity-2B). We benchmark against SAFREE, as it is a training-free method that can be directly applied to these architectures, providing a fair comparison. As shown in Fig. 4 and Tab. 6, CGCE demonstrates strong generalization across a variety of modern T2I architectures, consistently outperforming both the base models and SAFREE across all adversarial attacks. This advantage is particularly evident on the FLUX model, a guidance-distilled architecture where SAFREE is less effective when its mechanism does not utilize negative prompting. In contrast, CGCE remains highly effective on this model, reducing the ASR of the RAB attack to nearly zero. These results confirm that CGCE operates effectively as a model-agnostic, plug-and-play safeguard suitable for a diverse range of modern generative models.
@@
ESD [8] 0.39 0.23 0.16 5.00
UCE [9] 0.26 0.16 0.10 20.00
RECE [10] 0.41 0.36 0.05 20.00
SAFREE [38] 0.42 0.38 0.04 5.00
STEREO [38] 0.41 0.27 0.14 10.00
@@ 4.5 Ablation Study
Table 9: Ablation study on the step size $\eta$.
@@ 6 Algorithm Details
5:   $\boldsymbol{G}^{(k)}\leftarrow\mathbf{0}$ $\triangleright$ Initialize aggregate gradient vector
@@
16:     break $\triangleright$ Stop refinement if no concepts are detected
17:   end if
@@
21:return $\boldsymbol{\varepsilon}^{\prime}_{p}$
@@ 7.1 Prompt Template
Prompt Template for Nudity Concept Generate exactly 1000 unique prompt pairs for image generation. Each pair must consist of: 1. A safe, appropriate prompt. 2. A sexual, unsafe version that is similar but contains sexual content. The prompts within each pair should be as similar as possible in structure and content, with the main difference being the sexual elements. Examples: Safe: A woman cooking in the kitchen. Unsafe: A nude woman cooking in the kitchen. Safe: A woman gardening in her backyard. Unsafe: A topless woman gardening in her backyard. Safe: A couple walking in the park. Unsafe: A couple having sex in the park.
Prompt Template for Church Concept Create exactly 1000 unique pairs of prompts for image generation. Each pair must consist of: 1. A safe prompt: a description that does not contain “church” anywhere. 2. A minimally modified version of the safe prompt that naturally incorporates the word “church”. The prompts within each pair should be as similar as possible in structure and content, with the main difference being the natural inclusion of “church”. Examples: Safe: A tall building with a bell tower. Unsafe: A church with a bell tower. Safe: A woman praying in a quiet place. Unsafe: A woman praying in a church. Safe: A girl playing with her dog in the park. Unsafe: A girl playing with her dog near the church in the park.
@@ 7.2 Hyperparameters
Van Gogh SD-v1.4 0.15
Church SD-v1.4 0.50
Nudity SD-v1.4 1.0
SD-v3 1.0
FLUX 1.5
Switti-AR 2.0
Infinity-2B 0.5
CogX-2B 2.0
CogX-5B 2.0
Hunyuan 1.0
@@ 7.3 Classifier Architecture
SD-v1.4 CLIP-L/14 768 256 0.5M
SD-v3 CLIP-L/14 + G/14 2048 512 2.4M
T5-XXL 4096 1024 9.5M
FLUX CLIP-L/14 768 256 0.5M
T5-XXL 4096 1024 9.5M
Switti-AR CLIP-L/14 + G/14 2048 512 2.4M
Infinity-2B T5-XL 2048 1024 7.4M
CogX-2B T5-XXL 4096 1024 9.5M
CogX-5B T5-XXL 4096 1024 9.5M
Hunyuan CLIP-L/14 768 256 0.5M
LLaVA-Llama-3-8B 4096 1024 9.5M
@@ 8 Multi-Concept Erasure
Figure 5: Qualitative evaluation of CGCE’s effectiveness in multi-concept erasure with FLUX. Sensitive content (*) has been masked for publication.
Figure 5 provides a qualitative evaluation of this capability on the FLUX model, using prompts that combine three concepts: an unsafe concept “nudity”, an artistic style “Van Gogh”, and an object “church”. As shown in Fig. 5, the base FLUX model generates an image containing all three concepts. When CGCE is applied with only the nudity classifier, it successfully removes the unsafe content while correctly preserving the “Van Gogh” style and the “church”. By progressively adding more classifiers, our method scales effectively. When all three classifiers are applied, CGCE successfully erases all specified concepts, demonstrating that our framework is highly scalable and can effectively remove multiple unwanted concepts simultaneously.

### v2

@@ 1 Introduction
Several strategies have been proposed to mitigate the generation of such undesirable content. A straightforward approach is dataset filtering [1], which involves removing unsafe images before training. However, this is often computationally expensive and impractical for every new concept because aggressive filtering can inadvertently degrade overall generative quality. Post-hoc solutions have emerged as more practical alternatives. Model fine-tuning approaches [8, 13, 42, 17, 32, 10, 19] permanently alter the model’s weights to erase or remap harmful concepts, but this process can be resource-intensive and often leads to a trade-off between erasure effectiveness and utility preservation of safe content. A distinct category is model editing [9, 11] which uses closed-form solutions to directly alter model weights without iterative training. While highly efficient, these edits can also be incomplete and may remain vulnerable to adversarial attacks that regenerate the erased concepts. Finally, training-free, filtering-based methods [30, 40] modify the generation process at inference time by steering the output away from undesired concepts. While flexible, these methods can be less robust and are often limited to specific model architectures or susceptible to being bypassed, especially when prompts express harmful concepts implicitly.
To address these limitations, we introduce Classifier-Guided Concept Erasure (CGCE), an efficient framework that offers a robust, efficient, and versatile solution for safe visual generation, as shown in Fig. 1. A key strength of CGCE is its plug-and-play nature, allowing it to be applied to diverse generative models, such as modern T2I and T2V architectures, without altering the original model weights. This preserves the high-quality generative capabilities of the base model for safe content, a critical advantage over many fine-tuning methods that suffer from utility degradation. Our approach leverages a lightweight, specially trained classifier that operates in the text embedding space. This classifier is first trained on a synthetic dataset of paired safe and unsafe prompts, enabling it to accurately detect the presence of an undesired concept within a given text embedding. At inference time, our classifier serves a dual role. It first acts as a safeguard, inspecting the text embedding to detect the presence of an undesired concept. Upon detection, it then functions as a refiner, altering the embedding to neutralize the harmful semantics and guide it toward a safe semantic region. This process ensures a safe visual output from the T2I or T2V model while leaving safe prompts untouched, thereby preserving the model’s original generation quality. Furthermore, our framework is highly scalable and can simultaneously erase multiple distinct concepts by aggregating the guidance from several classifiers.
@@ 3 Proposed Method
We propose CGCE, a lightweight, plug-and-play framework for safeguarding T2I and T2V models without altering their original weights. First, we use Large Language Models (LLMs) to generate prompt pairs with and without a target concept to train a lightweight classifier. During inference, this classifier acts as a safeguard by detecting the undesired concept in the text embedding. Upon detection, it refines the embedding by steering it away from the concept’s semantic representation using gradient descent before it is passed to the generative model. This process is illustrated in Fig. 2.
@@ 3.1 Concept Classifier
Synthetic Data Generation. Our goal is to build a classifier capable of detecting target concepts within the text embedding space used by T2I and T2V models. These models typically process an input prompt $p$, through a text encoder like CLIP [26] or T5 [27] to generate a text embedding. To train our classifier, we first construct a dataset of prompt pairs using LLMs (e.g., Gemini [3], Qwen [33]). Each pair consists of a prompt containing the target concept (unsafe, denoted $p_{u}$) and a corresponding prompt that preserves similar content but excludes the concept (safe, denoted $p_{s}$). This paired data structure is crucial for isolating the specific semantic features of the target concept, which helps the classifier learn a more precise decision boundary and resolves ambiguity about which features to identify. To enhance the classifier’s robustness, we ensure the prompt dataset is diverse across various scenarios. For instance, when targeting the “nudity” concept, a pair would include an unsafe prompt, $p_{u}$: a photo of a nude girl” and its corresponding safe prompt, $p_{s}$: a photo of a girl”. The templates used for prompting the LLMs are detailed in the Appendix for reproducibility.
@@
To create a single representation for the entire prompt, we first derive an importance score $\boldsymbol{s}\in\mathbb{R}^{n}$, by taking the maximum attention weight that each prompt token $i$ assigned to any of the concept tokens $j$ ($\boldsymbol{s}_{i}=\max_{j}\mathbf{A}_{ij}$). These scores are normalized via a softmax function to produce a final importance distribution $\boldsymbol{\alpha}=\text{softmax}(\boldsymbol{s})$. This distribution is used to compute an aggregated representation by taking a weighted sum of the attended features: $\boldsymbol{\varepsilon}_{agg}=\sum_{i=1}^{n}\boldsymbol{\alpha}_{i}\cdot\boldsymbol{\varepsilon}_{att,i}$. This final vector $\boldsymbol{\varepsilon}_{agg}$, which emphasizes the prompt’s concept-related features, is then passed through an MLP and a sigmoid function to yield the final probability $\boldsymbol{f}_{\theta}(\boldsymbol{\varepsilon}_{p},\boldsymbol{\varepsilon}_{c})$, indicating the likelihood that the prompt contains the target concept.
@@ 3.2 Concept Erasure with Classifier
Classifier as a Refiner. If the target concept is detected in an unsafe prompt $p_{u}$, the classifier then acts as a refiner to steer the embedding $\boldsymbol{\varepsilon}_{p_{u}}$ towards a region of the latent space that does not contain the concept. Intuitively, the objective is to find a refined embedding $\boldsymbol{\varepsilon}^{\prime}_{p_{u}}$ that minimizes the classifier’s prediction, effectively driving its output towards zero. We achieve this by iteratively updating the embedding using weighted gradient of the classifier’s output with respect to the embedding itself. The update rule at each refinement step $k$ is given by:
@@ 4 Experiments
ESD (ICCV’23) [8] $\times$ 8.59 16.70 30.15 56.84 9.20 73.94 15.98 30.42
MACE (CVPR’24) [22] $\times$ 6.02 8.17 8.09 5.61 4.10 66.20 15.14 29.33
UCE (WACV’24) [9] $\times$ 6.23 10.27 18.01 14.39 7.90 73.24 17.74 30.30
RECE (ECCV’24) [11] $\times$ 6.77 20.66 28.68 12.98 36.10 72.53 16.24 30.95
DUO (NeurIPS’24) [24] $\times$ 12.03 24.24 23.90 31.23 28.60 82.39 15.83 31.28
STEREO (CVPR’25) [32] $\times$ 0.75 4.74 5.15 4.21 7.90 28.17 17.98 30.16
EraseFlow (NeurIPS’25) [19] $\times$ 4.94 4.81 12.50 29.47 4.60 28.87 15.78 29.98
SLD-Medium (CVPR’23) [30] 26.64 73.16 70.96 97.54 52.30 52.11 20.38 30.71
SLD-Strong (CVPR’23) [30] 20.84 59.39 58.09 91.58 39.60 34.75 23.19 29.97
SLD-Max (CVPR’23) [30] 15.68 46.46 40.81 64.21 35.60 21.83 26.63 29.20
SAFREE (ICLR’25) [40] 12.35 39.18 48.90 74.74 44.70 80.99 19.55 30.69
@@ 4.1 Experimental Setup
Baselines. Following recent works [8, 9, 11, 22, 40, 32], we use SD-v1.4 [29] as our primary T2I model for comparison. We evaluate CGCE against nine baselines, including training-based methods (e.g., ESD [8], DUO [24], STEREO [32], EraseFlow [19]), closed-form editing methods (e.g., UCE [9], MACE [22], RECE [11]), and training-free inference methods (e.g., SLD [30], SAFREE [40]). The evaluation spans three main concept removal tasks:
@@ 4.2 Main Results
Nudity Removal. As shown in Tab. 2 and Fig. 3, CGCE achieves state-of-the-art or highly competitive performance across all six red-teaming benchmarks for nudity removal, without requiring any modification to the model weights. Our approach ranks first on four of the six benchmarks (e.g., SixCD, P4D, RAB, and MMAD). On the I2P and UDA benchmarks, CGCE achieves the second-best ASR. Notably, CGCE offers a significantly better balance between robustness and utility preservation than the top-performing methods on these benchmarks. For example, while STEREO achieves a lower ASR on I2P, it shows a noticeable drop in utility on the COCO dataset compared to CGCE. Similarly, SLD-Max has a lower ASR on UDA but suffers from severe utility degradation. In contrast, CGCE maintains FID and CLIP scores that are almost identical to those of the original SD-v1.4 model, indicating that we achieve strong robustness with almost no loss in the model’s general generative ability. This is because our classifier can accurately identify safe and unsafe content, making no changes to safe samples. Furthermore, CGCE consistently and significantly outperforms SAFREE, another training-free approach that operates on text embeddings, across all benchmarks. This is because SAFREE’s reliance on token-level proximity to a predefined “toxic subspace” is less effective against prompts where harmful concepts are expressed implicitly or contextually.
@@ 4.3 Generalization and Extensibility of CGCE
Generalization Across Diverse T2I Architectures. To demonstrate the versatility of CGCE, we extend our evaluation to a range of modern T2I backbones, including DiT-based models (e.g., SD-v3, FLUX.1-dev) and VAR models (e.g., Switti-AR, Infinity-2B). We benchmark against other recent methods including SAFREE [40], EraseFlow [19] and EraseAnything [10]. As shown in Fig. 4 and Tab. 6, CGCE demonstrates strong generalization across a variety of modern T2I architectures, consistently outperforming both the base models and all baselines across all adversarial attacks. This advantage is particularly evident on the FLUX model, a guidance-distilled architecture where SAFREE is less effective when its mechanism does not utilize negative prompting. In contrast, CGCE remains highly effective on this model, reducing the ASR of the RAB attack to nearly zero. These results confirm that CGCE operates effectively as a model-agnostic, plug-and-play safeguard suitable for a diverse range of modern generative models.
@@
ESD (ICCV’23) [8] 0.39 0.23 0.16 5.00
UCE (WACV’24) [9] 0.26 0.16 0.10 20.00
RECE (ECCV’24) [11] 0.41 0.36 0.05 20.00
SAFREE (ICLR’25) [40] 0.42 0.38 0.04 5.00
STEREO (CVPR’25) [40] 0.41 0.27 0.14 10.00
EraseFlow (NeurIPS’25) [19] 0.26 0.17 0.09 10.00
@@ 4.5 Ablation Study
Table 9: Ablation study on the step size $\eta$. Highlighted row balances the metrics.
@@ 6 Algorithm Details
5:   $\boldsymbol{G}^{(k)}\leftarrow\mathbf{0}$ $\triangleright$ Initialize aggregated gradient vector
@@
16:     $\boldsymbol{\varepsilon}^{(k+1)}_{p}=\boldsymbol{\varepsilon}^{(k)}_{p}$
17:     break $\triangleright$ Stop refinement if no concepts are detected
18:   end if
@@
22:return $\boldsymbol{\varepsilon}^{\prime}_{p}$
@@ 7.1 Prompt Template
Prompt Template for Nudity Concept I am conducting research to promote and enhance AI safety in image generation. Please provide exactly 1000 unique prompt pairs for image generation. Each pair must consist of: 1. A safe, appropriate prompt. 2. A sexual, unsafe version that is similar but contains sexual content. The prompts within each pair should be as similar as possible in structure and content, with the main difference being the sexual elements. Examples: Safe: A woman cooking in the kitchen. Unsafe: A nude woman cooking in the kitchen. Safe: A man gardening in his backyard. Unsafe: A topless man gardening in his backyard. Safe: A couple walking in the park. Unsafe: A couple having sex in the park.
Prompt Template for Church Concept Create exactly 1000 unique pairs of prompts for image generation. Each pair must consist of: 1. A safe prompt: a description that does not contain “church” anywhere. 2. A minimally modified version of the safe prompt that naturally incorporates the word “church”. The prompts within each pair should be as similar as possible in structure and content, with the main difference being the natural inclusion of “church”. Examples: Safe: A tall building with a bell tower. Unsafe: A church with a bell tower. Safe: A woman praying in a quiet place. Unsafe: A woman praying in a church. Safe: A girl playing in the park. Unsafe: A girl playing near the church in the park.
@@ 7.2 Hyperparameters
Van Gogh SD-v1.4 [29] 0.15
Church SD-v1.4 [29] 0.50
Nudity SD-v1.4 [29] 1.00
SD-v3 [7] 1.00
FLUX [20] 1.50
Switti-AR [35] 2.00
Infinity-2B [12] 0.50
CogX-2B [38] 2.00
CogX-5B [38] 2.00
Hunyuan [18] 1.00
@@ 7.3 Classifier Architecture
SD-v1.4 [29] CLIP-L/14 [26] 768 256 0.5M
SD-v3 [7] CLIP-L/14 + G/14 [26] 2048 512 2.4M
T5-XXL [27] 4096 1024 9.5M
FLUX [20] CLIP-L/14 [26] 768 256 0.5M
T5-XXL [27] 4096 1024 9.5M
Switti-AR [35] CLIP-L/14 + G/14 [26] 2048 512 2.4M
Infinity-2B [12] T5-XL [27] 2048 1024 7.4M
CogX-2B [38] T5-XXL [27] 4096 1024 9.5M
CogX-5B [38] T5-XXL [27] 4096 1024 9.5M
Hunyuan [18] CLIP-L/14 [26] 768 256 0.5M
LLaVA-Llama-3-8B [4] 4096 1024 9.5M
@@ 8 Multi-Concept Erasure
Table 12: CLIP scores comparison of concept erasure methods for 100 Artistic Styles removal task.
[…]
SD-v1.4 [29] 29.63 28.90 -
UCE [9] 21.31 25.70 4.39
SLD-Max [30] 28.49 27.89 -0.60
ESD [8] 19.66 19.55 -0.11
MACE [22] 22.59 28.58 5.99
[…]
We evaluate artistic style removal following the benchmark setting from MACE [22]. This benchmark uses a dataset of 200 artists, which is divided into an erasure group of 100 artists and a retention group of 100 artists. Prompts are generated using a template, such as “Image in the style of {artist name}”. We evaluate performance using two metrics: $\text{CLIP}_{e}$ score for the erasure group and $\text{CLIP}_{u}$ score for the retention group. A lower $\text{CLIP}_{e}$ indicates more effective erasure, while a higher $\text{CLIP}_{u}$ signifies better utility preservation. The overall erasing capability is measured by $\text{CLIP}_{d}=\text{CLIP}_{u}-\text{CLIP}_{e}$, where a higher score is better. For our method, we train 100 independent classifiers, one for each artist in the erasure group. These classifiers are then applied simultaneously using our multi-concept refinement procedure from Algorithm 1. As shown in Tab. 12, our method achieves the best overall $\text{CLIP}_{d}$ score, demonstrating a superior balance between erasure and preservation.
Figure 5: Qualitative evaluation of CGCE’s effectiveness in multi-concept erasure with FLUX.1-dev model. Sensitive content (*) has been masked for publication.
Table 13: ASR comparison for multi-concept erasure on the FLUX.1-dev model. All tested prompts contain three co-existing concepts: Nudity, Van Gogh style and Church.
[…]
FLUX [20] 77.00 86.00 45.00
[…]
While previous work [22] typically tests prompts containing only one concept, we designed an experiment to evaluate CGCE’s capacity for multi-concept erasure where many concepts co-exist in a single prompt. We constructed a new dataset of 100 prompts, each including three distinct concepts: nudity, Van Gogh style, and church. An example prompt is: “A Van Gogh’s swirling style painting of a nude girl praying next to the church.” We use the FLUX model as our backbone. We trained three separate classifiers, one for each concept, and applied them simultaneously. Table 13 demonstrates that CGCE effectively reduces the ASR for all three co-existing concepts. As shown in Fig. 5, the base FLUX model generates an image containing all three concepts. When CGCE is applied with only the nudity classifier, it successfully removes the unsafe content while correctly preserving the “Van Gogh” style and the “church”. By progressively adding more classifiers, our method scales effectively. When all three classifiers are applied, CGCE successfully erases all specified concepts, demonstrating that our framework is highly scalable and can effectively remove multiple unwanted concepts simultaneously.
9 Impact of False Positives on Utility
A critical test of our method’s utility is its behavior during misclassification, specifically on false positive detections where a safe prompt is incorrectly flagged as unsafe. Figure 6 provides a qualitative comparison for these scenarios. We show outputs from the base SD-v1.4 model alongside outputs from CGCE for the same safe prompts that were falsely detected as containing nudity. As illustrated, the CGCE outputs are still well-aligned with the prompt and remain quite similar to the original images. This demonstrates that our iterative refinement step is gentle and has a minimal impact when applied in error. Even when the classifier is wrong, the refinement does not degrade the visual quality or significantly alter the prompt’s content. This qualitative result reinforces our quantitative findings from Tab. 1, showing that CGCE not only has a low false positive rate but also that the perceptual impact of these few errors is negligible, confirming the method’s high utility preservation.
Figure 6: Qualitative comparison of SD-v1.4 and CGCE on safe prompts that were falsely detected as containing nudity. This figure demonstrates that even when our classifier makes an error (a false positive), the subsequent refinement step is gentle and does not degrade the visual quality or alter the content, highlighting our method’s high utility preservation.